OpenStudy (kellanakins18):

2 squared plus 35 minus 32 plus 5 what is it

jhonyy9 (jhonyy9):

2^2 +35 -32 +5 = ?

jhonyy9 (jhonyy9):

this is your question please ?

OpenStudy (kellanakins18):

yes

jhonyy9 (jhonyy9):

ok . so step by step please collaborate

jhonyy9 (jhonyy9):

2^2 = ?

OpenStudy (kellanakins18):

4

jhonyy9 (jhonyy9):

correct 4+35 = ?

OpenStudy (kellanakins18):

39

jhonyy9 (jhonyy9):

correct 39-32 = ?

OpenStudy (kellanakins18):

7

jhonyy9 (jhonyy9):

right 7 +5 = ?

OpenStudy (kellanakins18):

12

jhonyy9 (jhonyy9):

exactly so this is the resulted right answer ok. ? hope you understand it now

OpenStudy (kellanakins18):

thank you i understand it now

jhonyy9 (jhonyy9):

welcome and was my pleasure good luck bye bye

OpenStudy (kellanakins18):

bye bye